# Enumeration of Elliptic Curves via Elliptic Gromov-Witten Invariants of Four Dimensional Projective Fano Hypersurfaces

### v4

@@ 1 Introduction
Now, let us turn to the case of genus $1$ GW invariants. The genus $1$ GW invariants of projective Fano surfaces (complex two-dimensional manifolds) are non-negative integers, at least for lower degrees. However, in the Calabi-Yau case, we know that the elliptic GW invariants include not only contributions from multiple cover maps from a genus $1$ curve but also degenerate contributions associated with genus 0 GW invariants [2, 15]. Roughly speaking, the degenerate contributions arise from the boundary loci of the moduli space of stable maps from genus $1$ stable curves, where the stable curves degenerate into nodal curves containing a genus 0 component. At these loci, the stable curves can be mapped into rational curves in the target manifold $X$, and the contributions are related to the genus 0 GW invariants of $X$. It is well-known that degenerate contributions also appear in the case of complex three-dimensional Fano manifolds. Heuristically, this phenomenon is expected from the dimensional counting of the virtual dimension of the moduli space of stable maps. The virtual dimension of the moduli space of stable maps from stable curves to complex three-dimensional projective manifolds with fixed degree $d$ is independent of the genus of the stable curves. Hence, we can speculate that the genus $1$ GW invariants may contain degenerate contributions because the operator insertions required to balance the positive virtual dimension of the moduli space do not depend on the genus. A natural question thus arises: **”Do degenerate contributions appear in the case of a projective Fano manifold whose complex dimension is greater than 3?”**
@@
Using these hints, we constructed a conjectural formula that describes the degenerate contributions in terms of genus 0 GW invariants, which will be presented in the next section. Heuristically, the genus $1$ GW invariants of degree $d$ are considered the sum of the degenerate contributions and the number of elliptic curves of the same degree satisfying the passing-through conditions. Therefore, if we subtract the degenerate contributions from the genus $1$ GW invariants, we expect to obtain non-negative integers. We denote the resulting number by $E_{d,a,b,c}$, and it turns out to be always a non-negative integer up to $d=5$. Specifically, $E_{d,a,b,c}=0$ if $d=1,2$, and it coincides with the number of elliptic curves computed by S. Katz’s method if $d=3$. The list of $E_{d,a,b,c}$ for $M_{6}^{k}$ ($k=1,2,3,4,5$) up to $d=5$ will be presented in Section 3. With these results, we propose a conjectural method for counting elliptic curves in a four-dimensional projective Fano hypersurface via genus $1$ GW invariants.
2 Definitions and Our Main Conjecture
@@
For later use, we also introduce the **Kähler (divisor) equation** and the **puncture equation** for the GW invariants.
$\langle{\cal O}_{h}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}=d\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}\quad(d\geq 1),$ (2.4)
@@
From (2.3), all $g=1$ and $d\geq 1$ GW invariants (given by (2.8)) vanish if $k>6$. In the $k=6$ case, the non-trivial genus 0 and $1$ GW invariants are given by $\langle{\cal O}_{h^{2}}\rangle_{0,d}$ and $\langle*\rangle_{1,d}$ ($d\geq 1$), respectively (* means no operator insertions). The relation between $\langle*\rangle_{1,d}$ and the number of elliptic curves of degree $d$ in $M_{6}^{6}$ was extensively studied in [14]. Therefore, we focus on the $k=1,2,3,4,5$ cases. In these cases, we evaluated the genus 0 GW invariants in (2.8) using both the multi-point virtual constants [10] and the associativity equation [13]. We then evaluated the corresponding genus $1$ invariants using the method of elliptic virtual structure constants [7]. Based on these numerical results and the three criteria introduced in Section 1, we propose the following conjecture:
@@
Heuristically, we expect that $E_{d,a,b,c}$ is the number of elliptic curves of degree $d$ in $M^{k}_{6}$ passing through $a$ 33-planes, $b$ 22-planes, and $c$ 11-planes in $CP^{5}**$. In particular, for $d=1,2$, our numerical computation confirmed that $E_{d,a,b,c}$ always vanishes. Furthermore, for $d=3$, we confirmed that $E_{d,a,b,c}$ coincides with the number of elliptic curves satisfying the required passing-through conditions, as computed by applying the method presented in [11] by S. Katz. We will explain the outline of S. Katz’s method in Appendix A.
3 Lists of $g=0,1$ Gromov-Witten Invariants and $E_{d,a,b,c}$ for $M^{k}_{6}$ ($k=1,2,3,4,5$) up to $d=5$
In this section, we present the numerical results used to determine the formula (2.9). For simplicity, we denote the genus $g$, degree $d$ Gromov-Witten (GW) invariants of $M_{6}^{k}$, $\left<({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\right>_{g,d}$, by $N^{g}_{d,a,b,c}$:
$N^{g}_{d,a,b,c}:=\left<({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\right>_{g,d}.$
The genus 0 GW invariants were computed using the **associativity equation** and the **multi-point virtual structure constants**, while the genus $1$ invariants were computed following the method described in [7].

### v5

@@ 1.1 Our Motivation
Now, let us turn to the case of genus $1$ GW invariants. The genus $1$ GW invariants of projective Fano surfaces (complex two-dimensional manifolds) are non-negative integers, at least for lower degrees. However, in the Calabi-Yau case, we know that the elliptic GW invariants include not only contributions from multiple cover maps from a genus $1$ curve but also degenerate contributions associated with genus 0 GW invariants [2, 16]. Roughly speaking, the degenerate contributions arise from the boundary loci of the moduli space of stable maps from genus $1$ stable curves, where the stable curves degenerate into nodal curves containing a genus 0 component. At these loci, the stable curves can be mapped into rational curves in the target manifold $X$, and the contributions are related to the genus 0 GW invariants of $X$. It is well-known that degenerate contributions also appear in the case of complex three-dimensional Fano manifolds. Heuristically, this phenomenon is expected from the dimensional counting of the virtual dimension of the moduli space of stable maps. The virtual dimension of the moduli space of stable maps from stable curves to complex three-dimensional projective manifolds with fixed degree $d$ is independent of the genus of the stable curves. Hence, we can speculate that the genus $1$ GW invariants may contain degenerate contributions because the operator insertions required to balance the positive virtual dimension of the moduli space do not depend on the genus. A natural question thus arises: ”Do degenerate contributions appear in the case of a projective Fano manifold whose complex dimension is greater than 3 ?”
@@
Using these hints, we constructed a conjectural formula that describes the degenerate contributions in terms of genus 0 GW invariants, which will be presented in the next section. Heuristically, the genus $1$ GW invariants of degree $d$ are considered the sum of the degenerate contributions and the number of elliptic curves of the same degree satisfying the passing-through conditions. Therefore, if we subtract the degenerate contributions from the genus $1$ GW invariants, we expect to obtain non-negative integers. We denote the resulting number by $E_{d,a,b,c}$, and it turns out to be always a non-negative integer up to $d=5$. Specifically, $E_{d,a,b,c}=0$ if $d=1,2$, and it coincides with the number of elliptic curves computed by S. Katz’s method if $d=3$. With these results, we propose a conjectural method for counting elliptic curves in a four-dimensional projective Fano hypersurface via genus $1$ GW invariants.
1.2 Definitions and Our Main Conjecture
@@
For later use, we also introduce the Kähler (divisor) axiom and the puncture axiom of the GW invariants.
$\langle{\cal O}_{h}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}=d\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}\quad(d\geq 1),$
@@
From (1.3), all $g=1$ and $d\geq 1$ GW invariants (given by (1.7)) vanish if $k>6$. In the $k=6$ case, the non-trivial genus 0 and $1$ GW invariants are given by $\langle{\cal O}_{h^{2}}\rangle_{0,d}$ and $\langle*\rangle_{1,d}$ ($d\geq 1$), respectively (* means no operator insertions). The relation between $\langle*\rangle_{1,d}$ and the number of elliptic curves of degree $d$ in $M_{6}^{6}$ was extensively studied in [15]. Therefore, we focus on the $k=1,2,3,4,5$ cases. In these cases, we evaluated the genus 0 GW invariants in (1.7) using both the multi-point virtual constants [10] and the associativity equation [14]. We then evaluated the corresponding genus $1$ invariants using the method of elliptic virtual structure constants [7] [^2] Text copies of Mathematica programs for these computations are available in ReserchGate homepage of Ken Kuwata[13].. Based on these numerical results and the three criteria introduced in Section 1, we propose the following conjecture:
@@
Heuristically, we expect that $E_{d,a,b,c}$ is the number of elliptic curves of degree $d$ in $M^{k}_{6}$ passing through $a$ 33-planes, $b$ 22-planes, and $c$ 11-planes in $CP^{5}$. In particular, for $d=1,2$, our numerical computation confirmed that $E_{d,a,b,c}$ always vanishes. Furthermore, for $d=3$, we confirmed that $E_{d,a,b,c}$ coincides with the number of elliptic curves satisfying the required passing-through conditions, as computed by applying the method presented in [11] by S. Katz. We will explain the outline of S. Katz’s method in Appendix A.
2 Derivation of the Conjecture
2.1 The $d=1$ case
We begin with the $d=1$ case. In this setting, it is known that no elliptic curves exist within the hypersurface. Consequently, the Gromov-Witten invariant $\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}$ consists solely of the degenerate contribution related to the genus 0 and degree $1$ invariant $\langle({\cal O}_{h^{2}})^{e}({\cal O}_{h^{2}})^{f}({\cal O}_{h^{2}})^{g}\rangle_{0,1}$.
In evaluating this degenerate contribution, the incidence conditions arising from operator insertions are effective. However, the virtual dimension of the genus 0 moduli space exceeds that of the genus $1$ moduli space by $1$. We therefore speculate that the contribution can be expressed using the genus 0 Gromov-Witten invariant $\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}$. Let us examine the data for these invariants for $M_{6}^{1}$, as shown in the tables below  [^3] For simplicity, we denote the genus $g$, degree $d$ Gromov-Witten invariants of $M_{6}^{k}$,
$\left<({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\right>_{g,d}$, by $N^{g}_{d,a,b,c}$.
Table 1: $g=0,M_{6}^{1},d=1$
d (a,b,c) $N^{0}_{d,a,b,c}$
1 (0,0,2) 11
1 (1,1,1) 11
1 (3,0,1) 11
1 (0,3,0) 11
1 (2,2,0) 22
1 (4,1,0) 33
1 (6,0,0) 55
Table 2: $g=1,M_{6}^{1},d=1$
d (a,b,c) $N^{1}_{d,a,b,c}$
1 (0,1,1) $\frac{5}{24}$
1 (2,0,1) $\frac{5}{24}$
1 (1,2,0) $\frac{5}{12}$
1 (3,1,0) $\frac{5}{8}$
1 (5,0,0) $\frac{25}{24}$
From this data, we readily identify the following relation for $k=1$:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{5}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}.$ (2.10)
Similar relations were found for other values of $k$:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{3}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}\quad(k=2),$
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{3}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}\quad(k=3),$
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{5}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}\quad(k=4),$
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{9}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}\quad(k=5).$ (2.11)
To unify these results, we assumed the following ansatz:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{ak^{2}+bk+c}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}.$
Solving for the coefficients, we obtained the general relation for $M_{6}^{k}$:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,1}=-\frac{k^{2}-5k+9}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,1}.$ (2.12)
2.2 Partial Results for the $d=2$ and $d=3$ cases
Next, we considered the $d=2$ case. In this case, there are no elliptic curves in $M_{6}^{k}$. Therefore, we initially assumed a relation of the form:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,2}=-\frac{\alpha(k,2)}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,2}.$ (2.13)
However, for cases other than $k=2$, additional contributions appear to exist, and the above ansatz does not hold. For the specific case $k=2$, we identified the following relation:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,2}=-\frac{5}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,2}\quad(k=2).$ (2.14)
We then turned to the $d=3$ case. Here, there is a non-trivial contribution to $\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,3}$ from the number of elliptic curves, denoted by $E_{3,a,b,c}$. At the early stages of this study, a general method for evaluating $E_{3,a,b,c}$ for arbitrary $M_{6}^{k}$ had not yet been established. However, for $k=3$, it is possible to count $E_{3,a,b,c}$ directly because a degree $3$ elliptic curve in $M_{6}^{3}$ is always given by the intersection of a $2$-plane in $CP^{5}$ and $M_{6}^{3}$.
Let $H_{a,b,c}$ be the number of $2$-planes that pass through $a$ 33-planes, $b$ 22-planes, and $c$ 11-planes in $CP^{5}$. This can be readily evaluated using Schubert calculus on the Grassmannian $Gr(3,6)$. Then, $E_{3,a,b,c}$ for $M_{6}^{3}$ is given by:
$E_{3,a,b,c}=3^{a+b+c}H_{a,b,c}\quad(k=3).$ (2.15)
Using the data obtained through this method, we found that the following relation holds for the $k=3,d=3$ case:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{1,3}=-\frac{5}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{2}})^{b}({\cal O}_{h^{2}})^{c}\rangle_{0,3}+E_{3,a,b,c}\quad(k=3).$ (2.16)
2.3 Incorporating the Results of Klemm and Pandharipande
In the $d=2$ and $d=3$ cases, we had not yet determined the additional contributions expected to appear, except for the specific cases mentioned in the previous subsection. To address this, we consulted the work of Klemm and Pandharipande [15], which provides the solution for the four-dimensional Calabi-Yau hypersurface $M_{6}^{6}$.
In the $M_{6}^{6}$ case, the non-trivial genus 0 and genus $1$ Gromov-Witten invariants of degree $d$ are:
$\langle{\cal O}_{h^{2}}\rangle_{0,d},\quad\langle*\rangle_{1,d},$
where the symbol * denotes the absence of operator insertions. Since $M_{6}^{6}$ is a Calabi-Yau manifold with vanishing first Chern class, these are the only non-vanishing invariants for these genera. To isolate the contributions of rational curves from multiple cover maps, we define the enumerative invariants $n_{0,d}$–representing the number of rational curves of degree $d$ passing through a $3$-plane in $CP^{5}$–via the following relation:
$\sum_{d=1}^{\infty}\langle{\cal O}_{h^{2}}\rangle_{0,d}q^{d}=\sum_{d=1}^{\infty}n_{0,d}\sum_{s=1}^{\infty}\frac{q^{sd}}{s^{2}}.$ (2.17)
The result in [15] for the Calabi-Yau hypersurface $M_{6}^{6}$ is given by:
$\sum_{d=1}^{\infty}\langle 1\rangle_{1,d}q^{d}$ $=$ $\sum_{d=1}^{\infty}n_{1,d}\sum_{s=1}^{\infty}\frac{\sigma(s)q^{sd}}{s}+\frac{1}{24}\sum_{d=1}^{\infty}c_{2}(M^{6}_{6})n_{0,d}\log(1-q^{d})$ (2.18)
$-\frac{1}{24}\sum_{d_{1},d_{2}=1}^{\infty}m_{d_{1},d_{2}}\log(1-q^{d_{1}+d_{2}}),$
where $n_{1,d}$ is the number of elliptic curves of degree $d$ in $M_{6}^{6}$ (corresponding to $E_{d,a,b,c}$ in our notation), $c_{2}(M^{6}_{6})=15$ is the second Chern class coefficient, and $m_{d_{1},d_{2}}$ is a rational number defined by the following recursive rules:
  1. 1.
    $m_{d_{1},d_{2}}=m_{d_{2},d_{1}}$
  2. 2.
    $m_{d_{1},d_{2}}=0$ if $d_{1}\leq 0$ or $d_{2}\leq 0$.
  3. 3.
    If $d_{1}\neq d_{2}$:
    $m_{d_{1},d_{2}}=\frac{n_{0,d_{1}}n_{0,d_{2}}}{k}+m_{d_{1},d_{2}-d_{1}}+m_{d_{1}-d_{2},d_{2}}$
  4. 4.
    If $d_{1}=d_{2}=d$:
    $m_{d,d}=c_{2}(M_{6}^{6})n_{0,d}+\frac{n_{0,d}n_{0,d}}{k}-\sum_{d_{1}+d_{2}=d}m_{d_{1},d_{2}}$
Formula (2.18) is complex due to multiple cover contributions arising from the vanishing first Chern class. However, in the Fano case, these contributions can often be ignored or simplified. We therefore hypothesized that a simpler structure exists and constructed the following ansatz for the Fano hypersurfaces:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{1,d}=E_{d,a,b,c}-\frac{\alpha(k,d)}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{0,d}$
$+\frac{\beta(k,d)}{24}\sum_{\begin{subarray}{c}d_{1}+d_{2}=d\\ d_{1},d_{2}>0\end{subarray}}\sum_{\alpha=1}^{3}\sum_{l=0}^{a}\sum_{m=0}^{b}\sum_{n=0}^{c}\binom{a}{l}\binom{b}{m}\binom{c}{n}$
$\times\frac{1}{k}\langle({\cal O}_{h^{2}})^{l}({\cal O}_{h^{3}})^{m}({\cal O}_{h^{4}})^{n}{\cal O}_{h^{\alpha}}\rangle_{0,d_{1}}\langle{\cal O}_{h^{4-\alpha}}({\cal O}_{h^{2}})^{a-l}({\cal O}_{h^{3}})^{b-m}({\cal O}_{h^{4}})^{c-n}\rangle_{0,d_{2}}.$ (2.19)
The three terms on the right-hand side of (2.19) correspond to the three components in (2.18). The second term, involving $\alpha(k,d)$, was partially determined from the $d=1,2$ cases as:
$\alpha(k,1)=k^{2}-5k+9\quad(1\leq k\leq 6),\quad\alpha(2,2)=5,\quad\alpha(3,3)=5.$ (2.20)
For $M_{6}^{6}$, (2.18) implies $\alpha(6,1)=c_{2}(M_{6}^{6})$. Given that the first and second Chern classes of $M_{6}^{k}$ are (6k)h(6-k)h and $(k^{2}-6k+15)h^{2}$ respectively, we identified $\alpha(k,1)$ as the combination $(k^{2}-6k+15)-(6-k)$. Following the logic that $\alpha(6,d)=15$ implied by (2.18), we generalized the ansatz to:
$\alpha(k,d)=(k^{2}-6k+15)-\gamma(d)(6-k).$
Using $\alpha(2,2)=5$ and $\alpha(3,3)=5$, we found $\gamma(2)=1/2$ and $\gamma(3)=1/3$, leading to our final form:
$\alpha(k,d)=(k^{2}-6k+15)-\frac{6-k}{d}.$ (2.21)
The third term in (2.19) is motivated by the boundary of the moduli space corresponding to nodal curves. Applying the splitting axiom, we assumed the singularity corresponds to the sum over operators $\sum\frac{1}{k}|{\cal O}_{h^{\alpha}}\rangle\langle{\cal O}_{h^{4-\alpha}}|$. Numerically testing this for $d=2$ (where $E_{2,a,b,c}=0$), we found:
$\beta(1,2)=-1/4,\quad\beta(2,2)=0,\quad\beta(3,2)=1/4,\quad\beta(4,2)=1/2,$
which suggests $\beta(k,2)=(k-2)/4$. For $k=5$, an exceptional extra term:
$\frac{1}{16}a!b!c!\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{0,1},$ (2.22)
appeared, likely because the virtual dimensions of the genus $1$ degree $2$ and genus 0 degree $1$ moduli spaces coincide.
For $d\geq 3$, we required that all $E_{d,a,b,c}$ derived from the ansatz be non-negative integers. This led to the refined ansatz:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{1,d}=E_{d,a,b,c}-\frac{\alpha(k,d)}{24}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{0,d}$
$+\frac{1}{48d}\sum_{\begin{subarray}{c}0<d_{1},d_{2}\\ d_{1}+d_{2}=d\end{subarray}}\beta(k,(d_{1},d_{2}))\sum_{\alpha=1}^{3}\sum_{l=0}^{a}\sum_{m=0}^{b}\sum_{n=0}^{c}\binom{a}{l}\binom{b}{m}\binom{c}{n}$
$\times\frac{1}{k}\langle({\cal O}_{h^{2}})^{l}({\cal O}_{h^{3}})^{m}({\cal O}_{h^{4}})^{n}{\cal O}_{h^{\alpha}}\rangle_{0,d_{1}}\langle{\cal O}_{h^{4-\alpha}}({\cal O}_{h^{2}})^{a-l}({\cal O}_{h^{3}})^{b-m}({\cal O}_{h^{4}})^{c-n}\rangle_{0,d_{2}}\quad(d\geq 3),$
Our numerical experiments yielded the following values for $\beta(k,(d_{1},d_{2}))$:
Table 3: Numerical values for $\beta(k,(d_{1},d_{2}))$
$(d_{1},d_{2})$ $\beta(k,(d_{1},d_{2}))$ $(d_{1},d_{2})$ $\beta(k,(d_{1},d_{2}))$
(1,1) k2k-2 (1,4) 4k144k-14
(1,2) 2k62k-6 (2,2) 4k164k-16
(1,3) 3k103k-10 (2,3) 6k266k-26
This allowed us to generalize the coefficient as $\beta(k,(d_{1},d_{2}))=2d-d_{1}d_{2}(6-k)$, leading to the final conjecture. In the tables below, we present our numerical data and the resulting $E_{d,a,b,c}$’s in the case of $M_{6}^{5}$. Subsequent validation using the method of S. Katz (Appendix A) confirmed the $d=3$ case.
Table 4: $g=0,M_{6}^{5}$
d (a,b,c) $N^{0}_{d,a,b,c}$
1 (0,1,0) $3250$
1 (2,0,0) $6125$
2 (0,0,1) $247500$
2 (1,1,0) $3718750$
2 (3,0,0) $17406875$
3 (1,0,1) $659250000$
3 (0,2,0) $2700512500$
3 (2,1,0) $19190225000$
3 (4,0,0) $150549428125$
4 (0,1,1) $529823250000$
4 (2,0,1) $4729124250000$
4 (1,2,0) $20567866625000$
4 (3,1,0) $209913851312500$
4 (5,0,0) $2337181124531250$
5 (0,0,2) $109236016800000$
5 (1,1,1) $5092187634000000$
5 (3,0,1) $61252356251250000$
5 (0,3,0) $21811124012125000$
5 (2,2,0) $274577525136875000$
5 (4,1,0) $3688919538904687500$
5 (6,0,0) $53412041211701171875$
Table 5: $g=1,M_{6}^{5}$
d (a,b,c) $N^{1}_{d,a,b,c}$ $E_{d,a,b,c}$
1 (1,0,0) $\frac{18375}{8}$ 0
2 (0,1,0) $\frac{8038625}{6}$ 0
2 (2,0,0) $\frac{148120375}{24}$ 0
3 (0,0,1) 246708750-246708750 $947500$
3 (1,1,0) 7059310625-7059310625 $14139375$
3 (3,0,0) $\frac{1319375639375}{24}$ $93667500$
4 (1,0,1) 1762031831250-1762031831250 $34818450000$
4 (0,2,0) $\frac{23015099134375}{3}$ $111635590625$
4 (2,1,0) $\frac{155974421459375}{2}$ $1070162215625$
4 (4,0,0) $\frac{5194583816496875}{6}$ $10767522628125$
5 (0,1,1) 1870863418500000-1870863418500000 $97094656425000$
5 (2,0,1) 22353636563531250-22353636563531250 $1193292296250000$
5 (1,2,0) $\frac{303154762779953125}{3}$ $4376885918562500$
5 (3,1,0) $\frac{8130487814000265625}{6}$ $56589153313015625$
5 (5,0,0) $\frac{470853639798192671875}{24}$ $767616850277828125$
  • Case $k=1$: Since $M_{6}^{1}$ is biholomorphic to $CP^{4}$, the moduli space is $\pi_{P}:P(S^{3}U^{*})\rightarrow Gr(3,5)$.
  • Case $k=2$: For a degree $3$ elliptic curve in $CP^{5}$ to be contained in $M_{6}^{2}$, the $2$-plane that contains the elliptic curve must itself be contained in $M_{6}^{2}$. Hence, the moduli space is given by $\pi_{P}^{-1}(PD(c_{\text{top}}(S^{2}U^{*})))$, where $PD(\alpha)$ is the submanifold of $Gr(3,6)$ that is Poincaré dual to the cohomology class $\alpha\in H^{*}(Gr(3,6))$, and $\pi_{P}:P(S^{3}U^{*})\rightarrow Gr(3,6)$ is the projection of the fiber space.
  • Cases $k=3,4,5$: For a degree $3$ elliptic curve in $CP^{5}$ to be contained in $M_{6}^{k}$, the restriction of the defining equation of $M_{6}^{k}$ (a homogeneous degree $k$ polynomial in $6$ homogeneous coordinates of $CP^{5}$) to the $2$-plane containing the elliptic curve must be divisible by the defining equation of the elliptic curve. Let $\pi_{S}:S_{P}\rightarrow P(S^{3}U^{*})$ be the tautological line bundle of the projectivization $P(S^{3}U^{*})$. Using this line bundle and the aforementioned condition, the moduli space of degree $3$ elliptic curves in $M_{6}^{k}$ is given by $PD(c_{\text{top}}(S^{k}U^{*}/(S^{k-3}U^{*}\otimes S_{P})))$.
$\frac{1}{1-c_{1}t+c_{2}t^{2}-c_{3}t^{3}}=1+\sum_{i=1}^{\infty}h_{i}t^{i}.$ (A.24)
Then, we have
$H^{*}(Gr(3,N))\simeq{\bf C}[c_{1},c_{2},c_{3}]/(h_{N-2},h_{N-1},h_{N}),$ (A.25)
$R(z,c_{1},c_{2},c_{3}):=z^{10}+\sum_{j=1}^{10}c_{i}(S^{3}U^{*})z^{10-i}=0.$ (A.26)
Hence, the ring structure of $H^{*}(P(S^{3}U^{*}))$ is given by
$\int_{P(S^{3}U^{*})}z^{9}\cdot(c_{3})^{N-3}=1.$ (A.28)
$\pi_{F*}(ev^{*}(h^{2}))$ $=$ $3c_{1}+z=:G_{1}$
$\pi_{F*}(ev^{*}(h^{3}))$ $=$ $3c_{1}^{2}-3c_{2}+c_{1}z=:G_{2}$
$\pi_{F*}(ev^{*}(h^{4}))$ $=$ $3c_{1}^{3}-6c_{1}c_{2}+3c_{3}+(c_{1}^{2}-c_{2})z=:G_{3}.$
We find that $\tilde{E}_{3,a,b,c}$ indeed coincides with $E_{3,a,b,c}$  [^4] A Text copy of the Maple worksheet used for computing $\tilde{E}_{3,a,b,c}$ of $M_{6}^{5}$ is available on M. Jinzenji’s ResearchGate homepage [9].. In the $k=5$ case, we obtain the following results:
$\tilde{E}_{3,0,0,1}=947500,\quad\tilde{E}_{3,1,1,0}=14139375,\quad\tilde{E}_{3,3,0,0}=93667500.$ (A.29)
Appendix B Lists of $g=0,1$ Gromov-Witten Invariants and $E_{d,a,b,c}$ for $M^{k}_{6}$ ($k=1,2,3,4$) up to $d=5$